Use appropriate algebra and Theorem 7.2.1 to find the given inverse Laplace transform. (Write your answer as a function of t.) ℒ−1 1 s2 + s − 12
Answer #1

aiven inverse Laplace transform AUS Plis-is] partial Fractions are B 1 1 s+4 s²+5-12 (5+4) (33) རྒྱུ་ 2 AL33) + Bls+4) stuss}AABEO -BA+UB=1 solving and . Then 0x3 => ZA+B=0 Ixls) BA & 4B El 73 Ela D BE A AS BE ř , then As substitute As. - I Be Ž in D->f415 - 11 szü} + 2 El =flt)= e-4t Žest flt)= Ilest tt). la Answer flt) = Lest eht]